OpenStudy (anonymous):

find ∫G(x,y)ds ??? , on the indicated curve G(X,Y)=2XY, C: X=5*COST, Y=5*SINT, 0≤t≤π÷4

OpenStudy (akshay_budhkar):

ok i will give u a hint, u should solve it yourself :D

OpenStudy (anonymous):

ok no problem

OpenStudy (akshay_budhkar):

\[2 \sin(x) \cos(x)=\sin(2x)\]

OpenStudy (akshay_budhkar):

so your G(x,y) =

OpenStudy (akshay_budhkar):

\[2 \times 5 \times \cos(x) \times 5 \times \sin(x)\]

OpenStudy (akshay_budhkar):

= \[25 \sin (2x)\]

OpenStudy (akshay_budhkar):

you can intergerate that :D

OpenStudy (anonymous):

i know how to find \[\int\limits_{C}^{} G(x,y) dx \] and \[\int\limits_{C}^{} G(x,y) dy \] what i want is \[\int\limits_{C}^{} G(x,y) ds \]

OpenStudy (akshay_budhkar):

one min i need to get a better help 1 second

OpenStudy (anonymous):

i am thinking that i should add \[\int\limits_{C}^{} G(x,y) dx + \int\limits_{C}^{} G(x,y) dy \] but i am not sure

OpenStudy (akshay_budhkar):

@turing ds???

OpenStudy (anonymous):

what do you think ?

OpenStudy (turingtest):

I'm sorry I'm multitasking...

OpenStudy (anonymous):

this is how the question written exactly ... what does he mean by ds?

OpenStudy (akshay_budhkar):

i don't know that is why i called turing

OpenStudy (akshay_budhkar):

ds? are u sure that is what is asking for?

OpenStudy (anonymous):

yes i am sure

OpenStudy (anonymous):

the answer will be 125/2 .. but i do not know how to get it

OpenStudy (akshay_budhkar):

why do you think you need to add?

OpenStudy (akshay_budhkar):

ds generally in my questions is the area

OpenStudy (anonymous):

really i am not sure i tried everything but i could not reach the answer

OpenStudy (akshay_budhkar):

i am waiting for turing to respond lol.. he will surely help.. i m thinking of possibilities till then

OpenStudy (anonymous):

ok I appreciate your help bro .. thanks.

OpenStudy (akshay_budhkar):

no no its not done! i wont sleep till u get ur answer :P

OpenStudy (anonymous):

Lol thanks so much.. me too cuz i have to submit this assignment tomorrow :D

OpenStudy (turingtest):

ok ds=sqrt[(dx/dt)^2+(dy/dt)^2]=5 2xy=50cost*sint got it from there?

OpenStudy (anonymous):

Integrating with respect to ds means that you are integrating with respect to the arc length (which is often denoted s) of the curve C. Note that \[ds=\sqrt{dx^2+dy^2}=\sqrt{(-5\cos (t) dt)^2+(5\sin (t) dt)^2}=5dt\] Thus, you have that \[\int_C G(x,y)ds=\int_0^{\pi/4}25\sin(2t)\cdot 5dt = \int_0^{\pi/4}125\sin(2t)dt = \frac{125}{2}\]

OpenStudy (akshay_budhkar):

ds?

OpenStudy (akshay_budhkar):

arc length is what it signifies?

OpenStudy (turingtest):

yes, you need ds in line integrals, arc length differential

OpenStudy (anonymous):

thanks guys I get it .. I appreciate your help .